# SpikeStereoNet: A Brain-Inspired Framework for Stereo Depth Estimation from Spike Streams

### v1

@@ A.3 Details of the Real Dataset
The real dataset comprises some video sequences captured under diverse environmental conditions, including varying ambient illumination levels (e.g., low, medium, and high brightness) and a range of dynamic scenarios characterized by distinct object arrangements. These scenes encompass different spatial configurations and relative motions of multiple objects (e.g., camera shake and object motion), introducing varied degrees of motion blur. Each blurry input frame is paired with 100 corresponding spike frames, forming a substantial collection of blur-spike pairs suitable for training and evaluation.
@@
Figure A3: The photograph of our hybrid imaging system. The RGB-D camera is fixed on the top plane of the stereo spike cameras. The extrinsic parameters between the cameras are calibrated.
A.4 More Experimental Results
A.4.1 Additional Comparative Results on the Synthetic Dataset
@@
We present a complete stereo depth estimation sequence for a single scene within the synthetic dataset, capturing the full motion trajectory of objects throughout the dynamic environment. As illustrated in Fig. A4, the predicted depth maps effectively reconstruct the continuous evolution of the scene, accurately reflecting both spatial structure and temporal consistency. These results demonstrate the model’s ability to preserve coherent geometry across time, validating its effectiveness in handling dynamic visual inputs. The supplementary materials in the zip file include continuously changing videos.
@@
Figure A4: Video results of our method on the synthetic dataset in the single scene. From left to right, the figure shows the sequential scene images, the visualized spike stream using the TFI method, the ground-truth depth maps, and the depth estimation results produced by our proposed method.
@@
A.4.2 Comparison on the Real Dataset
In the main experimental section, we have presented the qualitative results of several models on the real-world dataset. Here, we further report quantitative evaluation metrics to complement visual analysis. Specifically, we adopt the average end-point error (AvgErr) and the bad pixel ratio as our evaluation criteria. The detailed results of different models are summarized in Table  A2.
Table A2: Quantitative results on the real dataset, and all methods use real spike streams as input.
[…]
RAFT-Stereo [29] 6.18 3.39 0.64
DLNR [31] 5.64 3.38 0.61
Selective-stereo [36] 5.50 3.43 0.58
@@ A.4.2 Comparison on the Real Dataset
As shown in the last row of Table  A2, we can see that our method also outperforms other methods with the smallest AvgErr on the real dataset. The results of the experiment on the real dataset demonstrate that our method can effectively handle real scenarios.
A.4.3 Additional Results for Ablation Study
We provide additional ablation results to demonstrate the effectiveness of our designed modules. We analyze the impact of iteration steps by varying the number of iterations T in our RSNN-based model. As shown in Table  A3, our method achieves competitive or superior performance with significantly fewer iterations compared to existing approaches. Unlike conventional GRUs or transformer-based update modules, RSNNs require fewer parameters and operations per iteration, thus maintaining a lower computational cost.
Table A3: Ablation study for number of iterations and runtime.
@@ A.4.3 Additional Results for Ablation Study
Figure A5: Visualization results of ablation study for evaluating the effectiveness of our framework on the synthetic dataset. Among them, T = 32 is the baseline of our model.
In addition, to further support the ablation studies presented in the main experimental section, we provide the corresponding qualitative results in Fig.  A5. These visualizations clearly demonstrate the effectiveness of each module in our proposed framework and validate their individual contributions to the overall performance.
A.5 Broader Impacts
The proposed framework for stereo depth estimation from spike streams holds strong potential to advance neuromorphic vision in diverse application scenarios. In autonomous driving, the enhanced depth estimation enabled by spike-based sensing can improve 3D scene understanding, obstacle detection, and object localization, thereby contributing to more robust and reliable navigation systems. In the field of robotics, the fine-grained depth perception afforded by this method supports precise manipulation, mapping, and motion planning, especially in highly dynamic or high-speed environments where conventional cameras typically underperform. Moreover, this framework is the first that can directly estimate depth from raw stereo spike streams under supervised settings, enabling broader exploration of spike-based vision in both research and industrial applications.

### v2

@@ A.4.3 Real Dataset Collection
The real dataset comprises some video sequences captured under diverse environmental conditions, including varying ambient illumination levels (e.g., low, medium, and high brightness) and a range of dynamic scenarios characterized by distinct object arrangements. These scenes encompass different spatial configurations and relative motions of multiple objects (e.g., camera shake and object motion), introducing varied degrees of motion blur. The following is the specific collection process of our real dataset:
Dataset Acquisition Setup. Our real dataset is collected using a custom-built hardware system consisting of:
  • Two spike cameras arranged in a stereo configuration, with a baseline of 80 mm (calibrated for epipolar alignment).
  • A Kinect depth camera through hardware triggering, used to capture ground-truth depth for alignment.
Calibration of the Stereo Spike Camera System. Stereo spike cameras are calibrated following a rigorous protocol:
[…]
  • Both spike cameras captured synchronized spike streams of the target, which are converted to TFI (Temporal Filtered Integration) images with a 50 ms window to ensure clarity.
  • The calibration quality is validated by checking that the reprojection errors for the 3D chessboard corners are <1<1 pixel across all viewpoints.
Dataset Content. The dataset includes indoor real-world scenes:
  • Objects: 20+ everyday items, including textureless objects (e.g., plain plastic basin, white mugs) and reflective objects (e.g., metal box). Each scene contains 5-8 randomly arranged objects to simulate clutter.
  • Lighting Conditions: 8 controlled lighting setups with direct light sources (LED panels) at angles of $30^{\circ},45^{\circ},60^{\circ}$, and $90^{\circ}$ relative to the scene, with intensities ranging from 300-1500 lux to induce varying levels of specularity and shadow.
  • Data Modalities: For each scene, we provide raw stereo spike streams and Kinect-aligned depth maps (ground truth).
@@
Figure 10: The photograph of our hybrid imaging system. The RGB-D camera is fixed on the top plane of the stereo spike cameras. The extrinsic parameters between the cameras are calibrated.
A.5 More Experimental Results
A.5.1 Theoretical Analysis of Networks
We further supplement the theoretical analysis in the article with the explicit convergence of RSNN and Lipschitz stability analysis. Here is the complete proof:
Constructing Fixed Point Iteration and Mapping. For clarity, we rewrite the RSNN update step in vector form:
$\mathbf{u}_{t}=F(\mathbf{u}_{t-1})\quad\text{with}\quad\mathbf{u}:=(\mathbf{h},\mathbf{v},\mathbf{s},\mathbf{v}^{\text{th}})\in\mathbb{R}^{4N},$ (14)
$F(\mathbf{u})=\left[\begin{array}[]{l}\alpha_{t}\mathbf{v}+(1-\alpha_{t})\bigl(W_{\text{rec}}\theta(\mathbf{h}-\mathbf{v}^{\text{th}})+W_{f}\mathbf{s}^{(l-1)}\bigr)\\ \beta_{t}v_{\text{peak}}\\ \theta(\mathbf{h}-\mathbf{v}^{\text{th}})\\ \mathbf{h}-\gamma_{t}\theta(\mathbf{h}-\mathbf{v}^{\text{th}})\mathbf{v}^{\text{th}}\end{array}\right].$ (15)
The parameters $\bm{\alpha_{t}},\bm{\beta_{t}},\bm{\gamma_{t}}$ are all adaptive variables regulated by the firing rate, and after the activation function, their ranges are all (0, 1), that is, $0<\bm{\alpha_{t}}<1$, $0<\bm{\beta_{t}}<1$, $0<\bm{\gamma_{t}}<1$.
Lipschitz Continuity. The Lipschitz constant is determined by the activation function $\theta(\cdot)$ (hard threshold/ReLU) and a linear operator. If $\theta$ is 1-Lipschitz ReLU, the mapping $F$ is globally Lipschitz in $\mathbf{u}_{t}$, then:
$\|F(\mathbf{u})-F(\mathbf{u}^{\prime})\|\leq L\|\mathbf{u}-\mathbf{u}^{\prime}\|,$ (16)
where
$L=\alpha_{t}+(1-\alpha_{t})\,\gamma_{t}\,\beta_{t}\,v_{\text{peak}}\,\|W_{\text{rec}}\|,$ (17)
and all the ranges of variables in this equation are (0, 1), hence the RSNN update is a contraction with constant $L<1$.
Iterative Refinement and Convergence. Because $F$ is a contraction on the complete metric space $(\mathbb{R}^{N}\lVert\cdot\rVert)$, and according to the Banach Fixed-Point theorem (Gordji et al., 2017), the error satisfies.
$\|\mathbf{u}^{(k)}-\mathbf{u}^{*}\|\leq\frac{L^{k}}{1-L}\|\mathbf{u}^{(1)}-\mathbf{u}^{(0)}\|.$ (18)
Hence, the sequence $\{\mathbf{u}_{t}\}$ linear exponentially to a unique fixed point $\mathbf{u}^{\star}$. The bound agrees with the empirical spectral‑radius curves reported earlier (Fig. 7).
Practical Implications.
  • No Exploding States: The membrane potentials remain bounded by $\tfrac{1}{1-L}\|\mathbf{u}^{(0)}\|$.
  • Training Stability: Gradients passed through time have an upper bound $L^{T-t}$, which mitigates the exploding BPTT signals.
  • Noise Resilience: Any perturbation $\delta$ at step $k$ decays as $L^{\,t-k}\delta$.
Under the condition that $\bm{\alpha_{t}},\bm{\beta_{t}},\bm{\gamma_{t}}$ are bounded and the weight norm satisfies $L<1$, the iterative $\mathbf{u}_{t}=F(\mathbf{u}_{t-1})$ of RSNN linearly converges to a unique fixed point, providing a theoretical guarantee for continuous and contraction Lipschitz mapping. The above theoretical proof supplements the empirical stability study already proposed (Fig. 7).
A.5.2 Additional Comparative Results
@@
We present a complete stereo depth estimation sequence for a single scene within the synthetic dataset, capturing the full motion trajectory of objects throughout the dynamic environment. The different scene images below are obtained by rotating the stereo spike camera system from different views, and we have extracted some of the views at equal intervals for display. As illustrated in Fig. 11 and Fig. 12, the predicted depth maps effectively reconstruct the continuous evolution of the scene, accurately reflecting both spatial structure and temporal consistency. These results demonstrate the model’s ability to preserve coherent geometry across time, validating its effectiveness in handling dynamic visual inputs. The supplementary materials in the zip file include continuously changing videos.
@@
Figure 11: Video results of our method on the synthetic dataset in the single scene. From left to right, the figure shows the sequential scene images, the visualized spike stream using the TFI method, the ground-truth depth maps, and the depth estimation results produced by our proposed method.
@@
Figure 12: Visualization results of SpikeStereoNet on the synthetic dataset in a single scene. We presented depth estimation results of 12 views in the scene.
A.5.3 Comparison on the Real Dataset
In the main experimental section, we have presented the qualitative results of several models in the real-world dataset. Here, we also report quantitative evaluation metrics to complement visual analysis. Specifically, we adopt the average end-point error (AvgErr) and the bad pixel ratio as our evaluation criteria. The detailed results of different models are summarized in Table  4.
Table 4: Quantitative results on the real dataset, and all methods use real spike streams as input.
Method bad 2.0 $(\%)\downarrow$ bad 3.0 $(\%)\downarrow$ AvgErr $\downarrow$
RAFT-Stereo (Lipson et al., 2021) 6.18 3.39 0.64
DLNR (Zhao et al., 2023) 5.64 3.38 0.61
Selective-Stereo (Wang et al., 2024a) 5.50 3.43 0.58
@@ A.5.3 Comparison on the Real Dataset
As shown in the last row of Table  4, we can see that our method also outperforms other methods with the smallest AvgErr in the real dataset. The results of the experiment on the real dataset demonstrate that our method can effectively handle real scenarios.
A.5.4 Additional Results for Ablation Study
We provide additional ablation studies to demonstrate the effectiveness of our designed modules.
RSNN Module. Below we disentangle the impact of group normalization (GN) from that of our RSNN refinement by re‑running the key variants and reporting the results in the same metrics used in Table 5. The RNN is the same number of neurons as the original model.
These results demonstrate that both components, the RSNN architecture and the GN contribute to the performance, but neither alone is sufficient to achieve our full model’s results. Thus, the competitive performance of our framework arises from the synergy between the RSNN (optimized for spike-based temporal patterns) and the GN (stabilizing its dynamics).
Adaptation of Neuron Model. We conducted an explicit replacement experiment in which the adaptive LIF (ALIF) neurons used in our recurrent refinement were substituted with a classical LIF implementation that keeps all adaptive scalars fixed. In the following, we formalize the LIF classical, highlight the theoretical differences, and report quantitative results.
The classical LIF neuron, unlike our ALIF, uses fixed parameters (no adaptive variables $\bm{\alpha_{t}},\bm{\beta_{t}},\bm{\gamma_{t}}$). Its dynamics can be written as:
$\bm{h_{t}}$ $=\alpha\cdot\bm{v_{t-1}}+(1-\alpha)\cdot(W_{\text{rec}}\bm{s_{t-1}}+W_{f}\bm{s^{(l-1)}_{t})}$, (19)
$v^{\text{th}}$ $=\beta\cdot v_{\text{peak}},$
$\bm{s_{t}}$ $=\theta(\bm{h_{t}}-v^{\text{th}}),$
$\bm{v_{t}}$ $=\bm{h_{t}}-\gamma_{t}\cdot\bm{s_{t}}\cdot v^{\text{th}},$
where $\alpha$, $\beta$ and $\gamma$ are fixed constants (e.g., 0.5, 1.0, 1.0 in our baselines), not learnable or input dependent. This removes adaptivity to temporal patterns in spike streams. All other network components, learning rates, and losses were kept identical. We conducted an ablation study by replacing ALIF with classical LIF in the synthetic dataset (the same architecture, fixed $\alpha$, $\beta$ and $\gamma$). All other conditions of the experiment are the same, and the results are in Table 6.
From the above table, it can be seen that the overall performance decreases significantly after replacement with the LIF model. The ALIF is critical for our task because stereo spike streams require adaptive temporal processing. Adaptive gating markedly improves both stability and final accuracy with negligible overhead, justifying our choice of the novel ALIF over the classical LIF.
Number of Iterations. We analyze the impact of iteration steps by varying the number of iterations T in our RSNN-based model. As shown in Table 7, our method achieves competitive or superior performance with significantly fewer iterations compared to existing approaches. Unlike conventional GRUs or transformer-based update modules, RSNNs require fewer parameters and operations per iteration, thus maintaining a lower computational cost.
Figure 13: Visualization results of ablation study for evaluating the effectiveness of our framework on the synthetic dataset. Among them, T = 32 is the baseline of our model.
Table 5: Ablation study for RSNN and group normalization.
Method bad 2.0 $(\%)\downarrow$ bad 3.0 $(\%)\downarrow$ AvgErr (px) $\downarrow$
RNN 7.88 4.04 0.76
RNN + GN 7.28 3.41 0.66
RSNN 6.38 2.85 0.51
RSNN + GN (Ours) 4.13 2.38 0.42
Table 6: Ablation study for adaptation variables.
Method bad 2.0 $(\%)\downarrow$ bad 3.0 $(\%)\downarrow$ AvgErr (px) $\downarrow$
LIF (fixed $\alpha$, $\beta$, $\gamma$) 7.05 4.06 0.69
ALIF (Ours) 4.13 2.38 0.42
Table 7: Ablation study for number of iterations and runtime.
@@ A.5.4 Additional Results for Ablation Study
In addition, to further support the ablation studies presented in the main experimental section, we provide the corresponding qualitative results in Fig. 13. These visualizations clearly demonstrate the effectiveness of each module in our proposed framework and validate their individual contributions to the overall performance.
Image to Video. Regarding our model for generating continuous spike streams using the video frame interpolation method. Below we justify our design choice, quantify the interpolation error, and report a new control experiment based on Blender’s native sub‑frame rendering and optical flow supervision.
[…]
Table 8: Motion statistics in the synthetic datasets.
Quantile $50\%$ $75\%$ $90\%$ $95\%$
Motion (px) 0.31 0.57 1.12 1.89
From Table 8, more than $95\%$ of the frames exhibit sub-pixel motion, far below the motion threshold ($\approx 4$ px) at which EMA-VFI begins to introduce visible artifacts. Thus interpolation artifacts are negligible for the vast majority of the data.
After interpolation, frames are passed through the spike simulator. The conversion from interpolated frames to spike streams via the spike simulator further mitigates residual artifacts. The simulator introduces Poisson-like temporal noise and sparsity, which naturally smooths minor interpolation inconsistencies. Thus, even if an interpolated frame has minor inaccuracies, the temporal difference between two adjacent bins remains dominated by photon shot noise injected by the simulator. Regarding ground‑truth depth labels, which are extracted directly from Blender’s z‑buffer per original key‑frame, and they do not rely on the interpolated RGB images. During spike simulation, the disparity is linearly blended between key‑frames, guaranteeing label consistency regardless of interpolation artifacts.
To compare the effectiveness of the methods, we re-rendered full sequences with ground-truth optical flow from Blender, and regenerated a “SubFrame‑GT” variant. The ground‑truth optical flow exported via Blender’s Vector pass, generate image data and passed to the spike simulator. We re-train the model from scratch using the same hyper-parameters, and the results are as Table 9.
Table 9: Ablation study for Video frame interpolation methods.
Dataset variant bad 2.0 $(\%)\downarrow$ AvgErr (px) $\downarrow$
SubFrame‑GT (no interpolation) 0.44 4.43
Original (Ours) 0.44 4.52
The performance difference is $2\%$ / 0.1 pp, indicating that interpolation artifacts do not materially influence training or final accuracy. Overall, the two methods represent different patterns and approaches that can be chosen for specific tasks and datasets.
A.6 Broader Impacts
The proposed framework for stereo depth estimation from spike streams has a strong potential to advance neuromorphic vision in diverse application scenarios. In autonomous driving, the enhanced depth estimation enabled by spike-based sensing can improve 3D scene understanding, obstacle detection, and object location, thus contributing to more robust and reliable navigation systems. In the field of robotics, the fine-grained depth perception afforded by this method supports precise manipulation, mapping, and motion planning, especially in highly dynamic or high-speed environments where conventional cameras typically underperform. Moreover, this framework is the first that can directly estimate depth from raw stereo spike streams under supervised settings, enabling broader exploration of spike-based vision in both research and industrial applications.
Future Works. We will strengthen the benchmark with a broader coverage, and therefore we have launched a follow-up data collection campaign that will: (1) Expand the variety of the scene: indoor, outdoor and semi-outdoor environments with larger depth ranges and moving platforms; (2) Increase lighting diversity: day–night sweeps and high dynamic range sequences; (3) Enrich object categories: targeting $\geq$ 250 physical objects in 30 WordNet classes, including specular and transparent materials. The existing and extended dataset will be released publicly under a permissive license and the accompanying capture/annotation pipeline will be open‑sourced to facilitate community contributions.